# Bounding Causal Effects with Leaky Instruments

### v1

@@ Abstract
Instrumental variables (IVs) are a popular and powerful tool for estimating causal effects in the presence of unobserved confounding. However, classical approaches rely on strong assumptions such as the exclusion criterion, which states that instrumental effects must be entirely mediated by treatments. This assumption often fails in practice. When IV methods are improperly applied to data that do not meet the exclusion criterion, estimated causal effects may be badly biased. In this work, we propose a novel solution that provides partial identification in linear models given a set of leaky instruments, which are allowed to violate the exclusion criterion to some limited degree. We derive a convex optimization objective that provides provably sharp bounds on the average treatment effect under some common forms of information leakage, and implement inference procedures to quantify the uncertainty of resulting estimates. We demonstrate our method in a set of experiments with simulated data, where it performs favorably against the state of the art.
@@ 1 Introduction
Estimating causal effects from observational data can be challenging when treatments are not randomly assigned. While the task is doable under some structural assumptions [shpitser2008], most methods require access to data on potential confounders. This access cannot be generally guaranteed, since confounding variables may be unknown or difficult to measure. One common strategy for identifying causal effects under unobserved confounding relies on instrumental variables (IVs), which have a direct effect on the treatment but only an indirect effect on outcomes. For instance, single nucleotide polymorphisms (SNPs) often serve as IVs in genetic epidemiology, where they may be used to investigate the impact of phenotypes (e.g., cholesterol levels) on health outcomes (e.g., cancer) in Mendelian randomization studies [didelez2007].
@@
The IV model relies on three core conditions, formally defined in Sect. 2. Informally, we may describe IVs as variables that are (A1) relevant, i.e. associated with the treatment; (A2) unconfounded, i.e. independent of common causes between treatment and outcome; and (A3) exclusive, i.e. only affect outcomes through the treatment. Under some restrictions on structural equations, IVs can be used to recover causal effects despite the presence of unobserved confounding. Popular IV methods include two-stage least squares [angrist1995], as well as more recent nonlinear extensions based on deep learning [hartford2017deep, xu2021learning], kernel regression [singh2019kernel, muandet2020, sorowit2022], and generalized method of moments [bennet2019].
Of the three core conditions that characterize the IV setup, only (A1) can be immediately evaluated via observables. (A2) fails if the latent confounder between treatment and outcome is also a parent of some proposed IV(s). To continue with the Mendelian randomization example, this issue can arise when nearby variants are correlated, a phenomenon known as linkage disequilibrium [Reich2001]. (A3) fails if the proposed IV is a direct cause of the outcome (see Fig. 1). This can happen with complex traits in genetics, where one gene affects multiple seemingly unrelated systems through a process called horizontal pleiotropy [Solovieff2013]. If IV methods are naïvely applied in either case, resulting inferences can be severely biased [VanderWeele2014].
Since valid IVs may be impossible to identify a priori, several authors in recent years have proposed methods to estimate causal effects given just a set of candidate IVs (see Sect. 5). Details vary, but the goal is almost always to recover point estimates for the average treatment effect (ATE), possibly with associated confidence or credible intervals. We set a strictly more general target, relaxing (A3) to recover nontrivial bounds on this parameter, i.e. to partially identify the ATE. There is a long tradition of analytic and Bayesian methods for partial identification in IV models [Manski1990, chickering96, Balke1997], as well as more recent works that exploit the flexibility of stochastic gradient descent [kallus_zhou_2020, kilbertus2020class, hu2021]. Generally, a set of valid IVs is presumed—but some authors have considered the case where a single instrument is allowed to have a bounded effect on the outcome [ramsahai2012, Conley2012, silva2016].
@@
Figure 2: Causal diagram of the SEM described by Eqs. 1-5. Edge weights correspond to linear coefficients, while unobserved confounding effects are represented by the bidirected edge connecting $\epsilon_{x}$ and $\epsilon_{y}$. The dashed edge from $\bm{Z}$ to $Y$ denotes possible violations of (A3).
@@ 2 Problem Setup
where $\theta\in\mathbb{R}$ and $\bm{\beta},\bm{\gamma}\in\mathbb{R}^{d_{\bm{Z}}}$ are linear weights; and $\epsilon_{x},\epsilon_{y}\in\mathbb{R}$ are residuals with mean 00, standard deviation $\eta_{x},\eta_{y}\geq 0$, and correlation $\rho\in[-1,1]$. This latter parameter $\rho$ quantifies the magnitude and direction of unobserved confounding. To better interpret results, we assume all $Z$’s are on roughly the same scale (e.g., standardized to unit variance). We make no further assumptions about the distribution of $\bm{Z}$. Our goal is to bound $\theta$, which denotes the average treatment effect (ATE) of $X$ on $Y$.
In the classical nonparametric IV setting, we have a set of unobserved confounders $\bm{U}\in\mathbb{R}^{d_{\bm{U}}}$ with direct effects on both $X$ and $Y$. Then $\bm{Z}$ is a set of valid instruments if and only if the following conditions are satisfied: [^1] The “exogeneity” assumption in econometrics is sometimes equated with (A2), and sometimes with the conjunction of (A2) and (A3) (see, e.g., [wooldridge2009introductory, Ch. 15]). We avoid all talk of exogeneity to avoid confusion.
@@
Adapting these assumptions to a linear SEM, we posit that $\bm{U}$ is a parent of both noise variables satisfying $\epsilon_{x}\operatorname{\perp\!\!\!\perp}\epsilon_{y}\mid\bm{U}$ and equate (conditional) independence with (conditional) covariance of zero. Under these conditions, we may compute treatment effects via two-stage least squares (2SLS) [bt_iv]. For this procedure, we solve Eq. 1 with ordinary least squares (OLS) and substitute fitted values from this model for $X$ in Eq. 2, which is in turn solved via OLS. The resulting $\hat{\theta}^{\text{2SLS}}$ is our ATE estimate. Note that (A3) implies that $\lVert\bm{\gamma}\rVert=0$, since in this case each $Z_{j}\in\bm{Z}$ receives zero weight in Eq. 2.
@@
We call variables that satisfy (A1), (A2), and either form of $\tau$-exclusion leaky instruments. These features are technically observed confounders (at least those with nonzero $\gamma$ coefficients). Were it not for the unobserved confounding induced by $\rho$, causal effects could be calculated by integrating over $\bm{Z}$, as in the backdoor adjustment [pearl2009causality, Ch. 3.3]. Unfortunately, this option is unavailable when $\rho\neq 0$. By exploiting known leakage threshold(s), however, we show how to recover sharp bounds on the ATE.
@@ 3.1 Scalar $\tau$-exclusion
where $\bm{\Sigma}$ is the observational covariance matrix of $\{X,Y,\bm{Z}\}$ and $\bm{\Sigma}_{\mathcal{M}}$ is the model covariance matrix implied by Eqs. 1, 2 and 5.
@@
There exists a surjective function $g_{p}:\mathbb{R}\mapsto\mathbb{R}_{\geq 0}$ that maps values of the ATE $\theta$ to the $L_{p}$ norm of the leakage weights $\bm{\gamma}$:
@@
In the special case of $p=2$, the optimum is given by the standard OLS estimator $\check{\theta}_{2}=(\bm{\beta}\cdot\bm{\beta})^{-1}\bm{\beta}\cdot\bm{\alpha}$. Though closed form solutions are not available for arbitrary $p$—even in the well-studied case of $p=1$ [Pollard_1991, portnoy1997, chen_analysis_2008]—the value is easily computed via numerical methods. Note that $\check{\theta}_{p}$ is unique for any strictly convex $L_{p}$ norm, but may form a compact interval for $p\in\{1,\infty\}$.
@@
Define $h_{p}:=g_{p}\circ f$, such that $h_{p}:[-1,1]\mapsto\mathbb{R}_{\geq 0}$ maps vales of $\rho$ to $\lVert\bm{\gamma}\rVert_{p}$. For any $\check{\theta}_{p}$ (either a unique solution or any point on the compact interval of solutions), $h_{p}$ achieves its minimum at:
@@
Under the assumptions of Thm. 1, the exclusion criterion (A3) holds iff $\tau=\tau^{*}_{p}=\check{\tau}_{p}=0$, in which case $\theta^{*}=\check{\theta}_{2}=\theta^{\text{2SLS}}$.
@@
Analytic solutions are generally intractable for $p\neq 2$. However, Thm. 2 guarantees the existence and uniqueness of valid, sharp ATE bounds in the leaky IV model for any $p\geq 1$. These values can be readily computed with numerical methods, e.g. linear programming techniques [linear_opt]. For the $L_{2}$ case, we derive the following solution in closed form.
@@ 3.3 Inference
Thm. 2 provides an exact solution with the population covariance matrix $\bm{\Sigma}$. In practice, of course, all parameters must be estimated from finite data. We generally take the sample covariance matrix $\hat{\bm{\Sigma}}$ as our plug-in estimator, but many alternatives are possible. Numerous Bayesian [leonard1992, daniels1999, gelman_bayesian, Ch. 3.6] and penalized likelihood [Schafer2005, warton2008, won_condition-number-regularized_2012] methods have been proposed for this task, or the closely related task of estimating a regularized precision matrix [Friedman2007, cai2011, Mazumder2012]. Several of these options are implemented in our accompanying software package. These alternatives may be especially attractive in high-dimensional settings with a large number of leaky IVs to ensure a positive definite $\hat{\bm{\Sigma}}$.
We use a variety of estimators in our experiments below and augment the procedure with inference techniques. First, we describe a parametric test of the exclusion criterion, as foreshadowed by Corollary 1.1. In the linear IV model with $d_{\bm{Z}}\geq 2$, it is well known [kuroki2005, Chen_Tian_Pearl_2014, silva_shimizu_2017] that exclusion imposes a set of so-called tetrad constraints on covariance parameters of the form:
@@
We propose to test $H_{0}$ via Monte Carlo, estimating $\hat{\theta}^{\text{2SLS}}$ on the original data and creating a null covariance matrix $\bm{\Sigma}^{0}$ by replacing $\hat{\bm{\Sigma}}_{\bm{z}y}$ with $\bm{\Sigma}_{\bm{z}y}^{0}:=\hat{\bm{\Sigma}}_{\bm{z}x}\hat{\theta}^{\text{2SLS}}$. We assume that samples are distributed according to some $P_{\bm{\Sigma}}\in\mathcal{P}$, where $\mathcal{P}$ denotes a family of distributions parameterized by a covariance matrix $\bm{\Sigma}$—obvious examples include multivariate Gaussian and $t$-distributions, although alternatives such as multivariate binomial or Poisson distributions are also viable [krummenauer1998, jiang2021]. So long as we can sample data under fixed values of $\bm{\Sigma}$, we can perform the following test.
@@
Next, we introduce a nonparametric bootstrapping procedure [efron_bootstrap] to quantify the uncertainty of ATE bounds. Specifically, we draw $B$ many datasets of size $n$ by sampling with replacement from the input data and estimate the covariance matrix $\hat{\bm{\Sigma}}_{(b)}$ for each $b\in[B]$. Our target parameters are $(\theta^{-},\theta^{+})_{\tau,p}^{*}$, i.e. the bounds we would expect from a partial identification oracle (henceforth a $\bm{\Sigma}$-oracle) with knowledge of the population covariance matrix for observed variables $\{X,Y,\bm{Z}\}$. Note that even with access to the true $\bm{\Sigma}$, $\tau^{*}_{p}$ and $\theta^{*}$ remain unidentifiable, and so we distinguish between $\bm{\Sigma}$-oracles and oracles tout court, who are additionally omniscient with respect to latent parameters and therefore able to point identify the ATE. By contrast, with $\tau\in[\check{\tau}_{p},\tau^{*}_{p})$, a $\bm{\Sigma}$-oracle will produce invalid bounds that lie in the error region of Fig. 4 (see Thm. 1).
@@
We can smooth the bootstraps with kernel density estimation (see Sect. 4) or use a Bayesian bootstrap to get an approximate posterior distribution [rubin1981]. Either way, Thm. 4 provides a template for testing claims about whether, for instance, zero lies above or below the partial identification interval with high probability.
@@ 4 Experiments
For our benchmark experiments, we generate data according to Eqs. 1-5 with the following process:
@@
We present the mean and standard deviation of ATE estimates for a range of methods, computed across 50 runs of $n=1000$. For our own own approach, LeakyIV, we set $\tau=1.1\tau^{*}_{2}$ and shade the interval between our mean estimates for $(\hat{\theta}^{-},\hat{\theta}^{+})_{\tau,2}$. We benchmark against two classic methods—the backdoor adjustment and 2SLS—as an illustrative baseline. We also compare our results to two methods designed for causal inference with some invalid instruments: sisVIVE, which performs implicit feature selection via an $L_{1}$ penalty on the candidate IVs [kang2016]; and mode-based estimation (MBE), which treats invalid instruments as outliers that can be ignored using robust inference techniques [Hartwig2017]. We refer readers to the original papers for details on each.
@@
An alternative family of methods for modeling latent parameters in the IV setting is based on Bayesian inference [Shapland2019, bucur2020, Gkatzionis2021]. The goal in this approach is to estimate a posterior distribution for $\theta$, with partial identification bounds given by the upper and lower $\alpha$-quantiles of the credible interval. Rather than compare against some off the shelf method that does not explicitly encode our $\tau$-exclusion criterion, we design a Markov chain Monte Carlo (MCMC) sampler to model causal effects in the leaky IV setting (for details, see Appx. B.3). Due to the computational demands of MCMC sampling, we focus on the case where $d_{\bm{Z}}=5,\bm{\Sigma}_{\bm{zz}}$ is diagonal, and the SNR for both $X$ and $Y$ is 2, varying only $\rho$ across a range of six possible values.

### v2

@@ Abstract
Instrumental variables (IVs) are a popular and powerful tool for estimating causal effects in the presence of unobserved confounding. However, classical approaches rely on strong assumptions such as the exclusion criterion, which states that instrumental effects must be entirely mediated by treatments. This assumption often fails in practice. When IV methods are improperly applied to data that do not meet the exclusion criterion, estimated causal effects may be badly biased. In this work, we propose a novel solution that provides partial identification in linear systems given a set of leaky instruments, which are allowed to violate the exclusion criterion to some limited degree. We derive a convex optimization objective that provides provably sharp bounds on the average treatment effect under some common forms of information leakage, and implement inference procedures to quantify the uncertainty of resulting estimates. We demonstrate our method in a set of experiments with simulated data, where it performs favorably against the state of the art. An accompanying R package, leakyIV, is available from CRAN.
@@ 1 Introduction
Estimating causal effects from observational data can be challenging when treatments are not randomly assigned. While the task is doable under some structural assumptions [Rubin, 2005, Shpitser and Pearl, 2008, Pearl, 2009], most methods require access to data on potential confounders. This access cannot be generally guaranteed, since confounding variables may be unknown or difficult to measure. One common strategy for identifying causal effects under unobserved confounding relies on instrumental variables (IVs) [Wright, 1928, Bowden and Turkington, 1984, Angrist et al., 1996], which have a direct effect on the treatment but only an indirect effect on outcomes. For instance, single nucleotide polymorphisms (SNPs) often serve as IVs in genetic epidemiology, where they may be used to investigate the impact of phenotypes (e.g., cholesterol levels) on health outcomes (e.g., cancer) in Mendelian randomization studies [Smith and Ebrahim, 2004, Didelez and Sheehan, 2007, Lawlor et al., 2008].
@@
The IV model relies on three core conditions, formally defined in Sect. 2. Informally, we may describe IVs as variables that are (A1) relevant, i.e. associated with the treatment; (A2) unconfounded, i.e. independent of common causes between treatment and outcome; and (A3) exclusive, i.e. only affect outcomes through the treatment. Under some restrictions on structural equations, IVs can be used to recover causal effects despite the presence of unobserved confounding. Popular IV methods include two-stage least squares [Angrist and Imbens, 1995], as well as nonparametric extensions based on conditional moment restrictions [Newey and Powell, 2003, Newey, 2013, Bennett et al., 2019] and more recent works exploiting kernel regression [Singh et al., 2019, Muandet et al., 2020, Zhang et al., 2023] and neural networks [Hartford et al., 2017, Xu et al., 2021, Saengkyongam et al., 2022].
Of the three core conditions that characterize the IV setup, only (A1) can be immediately evaluated via observables. (A2) fails if the latent confounder between treatment and outcome is also a parent of some proposed IV(s). To continue with the Mendelian randomization example, this issue can arise when nearby variants are correlated, a phenomenon known as linkage disequilibrium [Reich et al., 2001]. (A3) fails if the proposed IV is a direct cause of the outcome (see Fig. 1). This can happen with complex traits in genetics, where one gene affects multiple seemingly unrelated systems through a process called horizontal pleiotropy [Solovieff et al., 2013]. If IV methods are naïvely applied in either case, resulting inferences can be severely biased [VanderWeele et al., 2014].
Since valid IVs may be impossible to identify a priori, several authors in recent years have proposed methods to estimate causal effects given just a set of candidate IVs (see Sect. 5). Details vary, but the goal is almost always to recover point estimates for the average treatment effect (ATE), possibly with associated confidence or credible intervals. We set a strictly more general target, relaxing (A3) to recover nontrivial bounds on this parameter, i.e. to partially identify the ATE. There is a long tradition of analytic and Bayesian methods for partial identification in IV models [Manski, 1990, Chickering and Pearl, 1996, Balke and Pearl, 1997], as well as more recent works that exploit the flexibility of stochastic gradient descent [Kallus and Zhou, 2020, Kilbertus et al., 2020, Hu et al., 2021]. Generally, a set of valid IVs is presumed—although some authors have considered the case where a single instrument is allowed to have a small effect on the outcome [Ramsahai, 2012, Conley et al., 2012, Silva and Evans, 2016].
@@
Figure 2: Causal diagram of the SEM described by Eqs. 1-3. Edge weights correspond to linear coefficients, while unobserved confounding effects are represented by the bidirected edge connecting $\epsilon_{x}$ and $\epsilon_{y}$. The dashed edge from $\bm{Z}$ to $Y$ denotes possible violations of (A3).
@@ 2 Problem Setup
where $\theta\in\mathbb{R}$ and $\bm{\beta},\bm{\gamma}\in\mathbb{R}^{d_{\bm{Z}}}$ are linear weights; and $\epsilon_{x},\epsilon_{y}\in\mathbb{R}$ are residuals with mean 00, standard deviations $\eta_{x},\eta_{y}\geq 0$, and correlation $\rho\in[-1,1]$. This latter parameter $\rho$ quantifies the magnitude and direction of unobserved confounding. To better interpret results, we assume all $Z$’s are on roughly the same scale (e.g., standardized to unit variance). We make no further assumptions about the distribution of $\bm{Z}$. Our goal is to bound $\theta$, which denotes the average treatment effect (ATE) of $X$ on $Y$.
In the classical nonparametric IV setting, we have a set of unobserved confounders $\bm{U}\in\mathbb{R}^{d_{\bm{U}}}$ with direct effects on both $X$ and $Y$. Then $\bm{Z}$ is a set of valid instruments if and only if the following conditions are satisfied: [^1] The “exogeneity” assumption in econometrics is sometimes equated with (A2), and sometimes with the conjunction of (A2) and (A3) (see, e.g., [Wooldridge, 2019, Ch. 15]). We avoid all talk of exogeneity to avoid confusion.
@@
Adapting these assumptions to a linear SEM, we posit that $\bm{U}$ is a parent of both noise variables satisfying $\epsilon_{x}\operatorname{\perp\!\!\!\perp}\epsilon_{y}\mid\bm{U}$ and equate (conditional) independence with (conditional) covariance of zero. Under these conditions, we may compute treatment effects via two-stage least squares (2SLS) [Bowden and Turkington, 1984]. For this procedure, we solve Eq. 1 with ordinary least squares (OLS) and substitute fitted values from this model for $X$ in Eq. 2, which is in turn solved via OLS. The resulting $\hat{\theta}^{\text{2SLS}}$ is our ATE estimate. Note that (A3) implies that $\lVert\bm{\gamma}\rVert=0$, since in this case each $Z_{j}\in\bm{Z}$ receives zero weight in Eq. 2.
@@
We call variables that satisfy (A1), (A2), and either form of $\tau$-exclusion leaky instruments. These features are technically observed confounders (at least those with nonzero $\gamma$ coefficients). Were it not for the unobserved confounding induced by $\rho$, causal effects could be calculated by integrating over $\bm{Z}$, as in the backdoor adjustment [Pearl, 2009, Ch. 3.3]. Unfortunately, this option is unavailable when $\rho\neq 0$. By exploiting known leakage threshold(s), however, we show how to recover sharp bounds on the ATE.
@@ 3.1 Scalar $\tau$-exclusion
where $\bm{\Sigma}$ is the observational covariance matrix of $\{X,Y,\bm{Z}\}$ and $\bm{\Sigma}_{\mathcal{M}}$ is the model covariance matrix implied by Eqs. 1, 2 and 3.
@@
There is a surjective function $g_{p}:\mathbb{R}\mapsto\mathbb{R}_{\geq 0}$ that maps values of the ATE $\theta$ to the $L_{p}$ norm of the leakage weights $\bm{\gamma}$:
@@
In the special case of $p=2$, the optimum is given by the standard OLS estimator $\check{\theta}_{2}=(\bm{\beta}\cdot\bm{\beta})^{-1}\bm{\beta}\cdot\bm{\alpha}$. Though closed form solutions are not available for arbitrary $p$—even in the well-studied case of $p=1$ [Pollard, 1991, Portnoy and Koenker, 1997, Chen et al., 2008]—the value is easily computed via numerical methods. Note that $\check{\theta}_{p}$ is unique for any strictly convex $L_{p}$ norm, but may form a compact interval for $p\in\{1,\infty\}$.
@@
Define $h_{p}:=g_{p}\circ f$, such that $h_{p}:[-1,1]\mapsto\mathbb{R}_{\geq 0}$ maps values of $\rho$ to $\lVert\bm{\gamma}\rVert_{p}$. For any $\check{\theta}_{p}$ (either a unique solution or any point on the compact interval of solutions), $h_{p}$ achieves its minimum at:
@@
Under the assumptions of Thm. 1, the exclusion criterion (A3) holds iff $\tau^{*}_{p}=\check{\tau}_{p}=0$, in which case $\theta^{*}=\check{\theta}_{2}=\theta^{\text{2SLS}}$.
@@
Analytic solutions are generally intractable for $p\neq 2$. However, Thm. 2 guarantees the existence and uniqueness of valid, sharp ATE bounds in the leaky IV model for any $p\geq 1$. These values can be readily computed with numerical methods, e.g. linear programming techniques [Bertsimas and Tsitsiklis, 1997]. For the $L_{2}$ case, we derive the following solution in closed form.
@@ 3.3 Inference
Thm. 2 provides an exact solution with the population covariance matrix $\bm{\Sigma}$. In practice, of course, all parameters must be estimated from finite data. We generally take the sample covariance matrix $\hat{\bm{\Sigma}}$ as our plug-in estimator, but many alternatives are possible. Numerous Bayesian [Leonard and Hsu, 1992, Daniels and Kass, 1999, Gelman et al., 2014, Ch. 3.6] and penalized likelihood [Schäfer and Strimmer, 2005, Warton, 2008, Won et al., 2012] methods have been proposed for this task, or the closely related task of estimating a regularized precision matrix [Friedman et al., 2007, Cai et al., 2011, Mazumder and Hastie, 2012]. Several of these options are implemented in our accompanying software package. These alternatives may be especially attractive in high-dimensional settings with a large number of leaky IVs to ensure a positive definite $\hat{\bm{\Sigma}}$.
We use a variety of estimators in our experiments below and augment the procedure with inference techniques. First, we describe a parametric test of the exclusion criterion, as foreshadowed by Corollary 1.1. In the linear IV model with $d_{\bm{Z}}\geq 2$, it is well known [Kuroki and Cai, 2005, Chen et al., 2014, Silva and Shimizu, 2017] that exclusion imposes a set of so-called tetrad constraints on covariance parameters of the form:
@@
We propose to test $H_{0}$ via Monte Carlo, estimating $\hat{\theta}^{\text{2SLS}}$ on the original data and creating a null covariance matrix $\bm{\Sigma}^{0}$ by replacing $\hat{\bm{\Sigma}}_{\bm{z}y}$ with $\bm{\Sigma}_{\bm{z}y}^{0}:=\hat{\bm{\Sigma}}_{\bm{z}x}\hat{\theta}^{\text{2SLS}}$. We assume that samples are distributed according to some $P_{\bm{\Sigma}}\in\mathcal{P}$, where $\mathcal{P}$ denotes a family of distributions parameterized by a covariance matrix $\bm{\Sigma}$—obvious examples include multivariate Gaussian and $t$-distributions, although alternatives such as multivariate binomial or Poisson distributions are also viable [Krummenauer, 1998, Jiang et al., 2021]. So long as we can sample data under fixed values of $\bm{\Sigma}$, we can perform the following test.
@@
Next, we introduce a nonparametric bootstrapping procedure [Efron, 1979] to quantify the uncertainty of ATE bounds. Specifically, we draw $B$ many datasets of size $n$ by sampling with replacement from the input data and estimate the covariance matrix $\hat{\bm{\Sigma}}_{(b)}$ for each $b\in[B]$. Our target parameters are $(\theta^{-},\theta^{+})_{\tau,p}^{*}$, i.e. the bounds we would expect from a partial identification oracle (henceforth a $\bm{\Sigma}$-oracle) with knowledge of the population covariance matrix for observed variables $\{X,Y,\bm{Z}\}$. Note that even with access to the true $\bm{\Sigma}$, $\tau^{*}_{p}$ and $\theta^{*}$ remain unidentifiable, and so we distinguish between $\bm{\Sigma}$-oracles and oracles tout court, who are additionally omniscient with respect to latent parameters and therefore able to point identify the ATE. By contrast, with $\tau\in[\check{\tau}_{p},\tau^{*}_{p})$, a $\bm{\Sigma}$-oracle will produce invalid bounds that lie in the error region of Fig. 4 (see Thm. 1).
@@
We can smooth the bootstraps with kernel density estimation (see Sect. 4) or use a Bayesian bootstrap to get an approximate posterior distribution [Rubin, 1981]. Either way, Thm. 4 provides a template for testing claims about whether, for instance, zero lies above or below the partial identification interval with high probability.
@@ 4 Experiments
For our benchmark experiments, we generate data according to Eqs. 1-3 with the following process:
@@
We present the mean and standard deviation of ATE estimates for a range of methods, computed across 50 runs of $n=1000$. For our own own approach, LeakyIV, we set $\tau=1.1\tau^{*}_{2}$ and shade the interval between our mean estimates for $(\hat{\theta}^{-},\hat{\theta}^{+})_{\tau,2}$. We benchmark against two classic methods—the backdoor adjustment and 2SLS—as an illustrative baseline. We also compare our results to two methods designed for causal inference with some invalid instruments: sisVIVE, which performs implicit feature selection via an $L_{1}$ penalty on the candidate IVs [Kang et al., 2016]; and mode-based estimation (MBE), which treats invalid instruments as outliers that can be ignored using robust inference techniques [Hartwig et al., 2017]. We refer readers to the original papers for details on each.
@@
An alternative family of methods for modeling latent parameters in the IV setting is based on Bayesian inference [Shapland et al., 2019, Bucur et al., 2020, Gkatzionis et al., 2021]. The goal in this approach is to estimate a posterior distribution for $\theta$, with partial identification bounds given by the upper and lower $\alpha$-quantiles of the credible interval. Rather than compare against some off the shelf method that does not explicitly encode our $\tau$-exclusion criterion, we design a Markov chain Monte Carlo (MCMC) sampler to model causal effects in the leaky IV setting (for details, see Appx. B.3). Due to the computational demands of MCMC sampling, we focus on the case where $d_{\bm{Z}}=5,\bm{\Sigma}_{\bm{zz}}$ is diagonal, and the SNR for both $X$ and $Y$ is 2, varying only $\rho$ across a range of six possible values.